OpenStudy (anonymous):

please help Solve the system of equations by the substitution method. If the system does not have a single ordered pair as a solution, state whether the system is inconsistent or dependent. x + 3y = 32 -3x + 2y = 3 A. (4, 10) B. No solution; inconsistent C. (5, 9) D. (-5, 10)

OpenStudy (anonymous):

x + 3y = 32 -3x + 2y = 3 solve either of them for y

OpenStudy (anonymous):

ok